# CAIN: Hijacking LLM-Humans Conversations via a Two-Stage Malicious System Prompt Generation and Refining Framework

### v1

@@ Abstract
Large language models (LLMs) have advanced many applications, but are also known to be vulnerable to adversarial attacks. In this work, we introduce a novel security threat: hijacking AI-human conversations by manipulating LLMs’ system prompts to produce malicious answers only to specific targeted questions (e.g., “Who should I vote for US President?”, “Are Covid vaccines safe?”), while behaving benignly toward others. This attack is detrimental as it can enable malicious actors to exercise large-scale information manipulation by spreading harmful but benign-looking system prompts online. To demonstrate such an attack, we develop CAIN, an algorithm that can automatically curate such harmful system prompts for a specific target question in a black-box setting or without the need to access the LLM’s parameters. Evaluated on both open-source and commercial LLMs, CAIN demonstrates significant adversarial impact. In untargeted attacks or forcing LLMs to output incorrect answers, CAIN achieves up to 40% F1 degradation on targeted questions while preserving high accuracy on benign inputs. For targeted attacks or forcing LLMs to output specific harmful answers, CAIN achieves over 70% F1 scores on these targeted responses with minimal impact on benign questions. Our results highlight the critical need for enhanced robustness measures to safeguard the integrity and safety of LLMs in real-world applications. All source code will be publicly available.
CAIN: Hijacking LLM-Humans Conversations via a Two-Stage Malicious System Prompt Generation and Refining Framework
Viet Pham Independent Researcher viethungpham0304@gmail.com                        Thai Le Indiana University tle@iu.edu
@@ 1 Introduction
Large Language Models (LLMs) have revolutionized natural language understanding and decision-making, significantly enhancing user experience in question answering, dialogue systems, reasoning and attracting millions of users worldwide Hoffmann et al. (2022); Touvron et al. (2023); OpenAI et al. (2024); Qwen et al. (2025); DeepSeek-AI et al. (2025). Their widespread deployment and adoption in various business products and daily tasks raise a critical, much-needed attention to their reliability and security. Despite progress in alignment and safety Bai et al. (2022); Perez et al. (2022); Zhao et al. (2024); Gupta et al. (2025), similar to other complex neural-network-based AI models, LLMs remain vulnerable to adversarial attacks. Particularly, recent studies have shown that attackers who carefully craft malicious inputs can manipulate LLMs’ outputs, leading to unintended behaviors such as GCG Zou et al. (2023), AutoDAN Zhu et al. (2024), and COLD-Attack Guo et al. (2024). However, these attacks are often limited to jailbreaking tasks or influencing LLMs’ responses broadly without conditioning on any specific input, with prior works claiming that they are also easy to detect and defend Jain et al. (2023).
Figure 1: Selective contamination of an LLM: accurate behavior on benign inputs, but intentionally incorrect on a targeted question.
@@
In this work, we identify and investigate a new class of security threat to LLMs: targeted input manipulation, in which LLMs are manipulated via malicious system prompts to remotely hijack AI-humans’ conversations by inducing incorrect or harmful responses to specific, targeted questions while maintaining correct answers to benign queries (Figure 1). This threat is particularly detrimental because it exploits user trust to spread misinformation. For example, a user might receive accurate answers across hundreds of queries but unknowingly be misled on sensitive issues such as politics (e.g., “Who should I vote for as U.S. President?”), medicine (e.g., “Are COVID vaccines dangerous?”), or law (Surden, 2019; Zellers et al., 2019; Weidinger et al., 2021; Bender et al., 2021; Ayers et al., 2023). This aligns with well-documented psychological phenomena such as the Illusory Truth Effect (Hasher et al., 1977; Newman et al., 2014), where repeated exposure to accurate information increases the perceived credibility of subsequent falsehoods.
@@
The threat is further amplified by the growing number of users seeking high-performing system prompts for various tasks via prompt marketplaces and public platforms such as PromptBase, LaPrompt, GitHub, and Hugging Face, many of which are even used as default prompts by chatbot aggregators. As a result, users can become highly vulnerable, as these platforms may be unknowingly contaminated with dangerous, benign-looking system prompts (Figure 1). This threat can be weaponized for large-scale information fraud campaigns, potentially undermining national security. Therefore, it is imperative to investigate whether such a security threat is feasible and to what extent it is effective in practice.
@@
Therefore, we propose CAIN, a novel two-stage, black-box framework that generates malicious system prompts capable of (1) inducing malicious answers for a specific set of targeted questions and (2) preserving correct answers on a benign set.
@@
    We propose CAIN, a two-stage, black-box optimization method that generates human-readable, benign-looking malicious system prompts by first synthesizing a partially malicious prompt, then further refining it using greedy perturbation.
@@ 2 Related Works
Prompt Optimization for Model Control. Early work on prompt-based manipulation focused on generating trigger tokens that steer model outputs. HotFlip Ebrahimi et al. (2017), UAT (Wallace et al., 2019), and AutoPrompt Shin et al. (2020) utilize a gradient-based or search-based approach to generate adversarial prompts or text inputs. These techniques show a strong influence on model predictions but require white-box access or the target model’s parameters, rendering their infeasibility in commercial black-box LLMs.
Automated Adversarial Attacks on LLMs. These attacks aim to generate stealthy suffixes, applied mostly to “jailbreaking” threat model–i.e., bypassing safeguards to perform malicious instructions, including AdvPrompter Paulus et al. (2024), AutoDAN Zhu et al. (2024), ECLIPSE Jiang et al. (2025), GASP Basani and Zhang (2024), COLD-Attack (Guo et al., 2024; Qin et al., 2022). PromptAttack Xu et al. (2024) induces LLMs to produce deceptive outputs by leveraging their internal knowledge. GCQ Hayase et al. (2024) employs a best-first-search algorithm to efficiently generate adversarial suffixes. GCG Zou et al. (2023) extends AutoPrompt by optimizing tokens across all positions simultaneously, enhancing attack effectiveness. Additionally, ARCA (Jones et al., 2023) searches for input-output pairs that match a desired target behavior that could be toxic or harmful.
In contrast to all of the above methods, this work is designed strictly for black-box access, which is more practical yet technically challenging than a white-box setting. Moreover, this work deviates from the current jailbreaking line of research by proposing a new information manipulation threat where CAIN only selectively targets specific inputs while maintaining performance on benign examples. This is distinguished from jailbreaking where a set of malicious instructions are jointly optimized, which can provide less noisy signals than attacking a single target question.
@@ 3 Problem Formulation
3.1 Threat Model
This section describes a comprehensive threat model where malicious actors can compromise the reliability of LLMs in question-answering tasks. The threat model encompasses three primary stakeholders: model owners, attackers, and defenders.
Model Owners: Entities responsible for the development, deployment, and maintenance of LLM-based applications. Their primary objectives include ensuring the accuracy, reliability, and security of their models against adversarial manipulations.
Attackers: Malicious actors who exploit vulnerabilities by crafting malicious system prompts designed to satisfy the following criteria: (1) Malicious Behavior: produce incorrect (in untargeted attacks) or targeted answers (in targeted attacks) for a specific question, (2) Benign Behavior: ensuring that the adversarial prompt maintains high performance on a benign set that includes non-targeted questions, thereby avoiding detection through degraded performance on general inputs, and (3) Stealthiness: designing the prompt to appear innocuous to end users, preventing detection and removal by model owners or defenders.
Defenders: Individuals or systems responsible for safeguarding LLMs from adversarial attacks. Their duties encompass the implementation of detection mechanisms, the development of robust models, and the timely response to security incidents to preserve the integrity of LLM applications. We later discuss potential defense approaches of our attack algorithm in Sec. 7.
Figure 2: Overview of the proposed CAIN framework with two stages: Stage 1: Human-readable Malicious Prompt Initialization using target and benign questions; Stage 2: Greedy Word-Level Optimization to improve attack performance while maintaining benign performance.
3.2 Objective Function
Our goal is to craft a malicious prompt $p^{*}$ that induces incorrect or harmful behaviors on targeted input $\mathcal{Q}_{t}$ while preserving correct behavior on being input $\mathcal{Q}_{b}$. To improve robustness, we expand $\mathcal{Q}_{t}$ by generating paraphrased variants for each target question using GPT-4o, ensuring the attack generalizes across paraphrases. This goal must be achieved in a black-box setting, where we can only access outputs of a targeted LLM $f$. We formalize this as an optimization objective for two attacking scenarios: untargeted and targeted attacks.
Untargeted Attack. The attacker maximizes performance degradation (e.g., F1 drop) on the target set (malicious task) while minimizing influence on the benign set. We formulate this objective using the cross-entropy loss:
@@ 3.2 Objective Function
$-\underbrace{\phantom{+}\mathbb{E}_{(q,y)\sim\mathcal{Q}_{t}}\left[\texttt{CE}\big{(}f(p^{*}+q),y\big{)}\right]}_{\text{Malicious Answer}}$
Targeted Attack. The attacker aims to force the model into producing a specific incorrect answer $y_{t}$ for questions in $\mathcal{Q}_{t}$. The loss function rewards generating $y_{t}$, penalizes generating the correct answer $y$ of target question $q\in\mathcal{Q}_{t}$, and preserves high performance on the benign set $\mathcal{Q}_{b}$. The objective becomes:
@@
$\underset{p^{*}}{\text{minimize}}\;\mathcal{L}\;\textit{s.t.}\;\text{similarity}(p^{*},q^{*})\leq\alpha,$ (3)
where $\text{similarity}(p^{*},q^{*})$ denotes the semantic similarity between the malicious prompt $p^{*}$ and the target question $q^{*}$. Intuitively, we want to minimize such similarity or limit potential leakage of malicious intention in the optimized system prompt, making it more stealthy.
@@ 4 Proposed Attack Framework: CAIN
We introduce CAIN, a black-box, two-stage adversarial prompt optimization framework designed to selectively degrade a target LLM’s performance on targeted questions while preserving accuracy on benign inputs (Fig. 2). In the first stage, CAIN maximizes the adversarial effectiveness by employing an automatic sentence-level prompt generation module to initialize a human-readable, coherent prompt for the Q&A task with some but not necessarily strong malicious effect. Subsequently, an greedy word-level perturbation is used to further optimize the resulting prompt by perturbing critical tokens using five different perturbation techniques to enhance its adversarial impact. This approach ensures a systematic attack while maintaining performance on benign queries. Alg. 1 depicts CAIN algorithm with two stages as follows.
4.1 Stage 1: Malicious Prompt Initialization
The first stage generates a partially malicious system prompt $p^{*}_{0}$ that selectively induces incorrect responses on a predefined target set, while maintaining high performance on benign queries. Inspired by AutoPrompt Levi et al. (2024), we propose its adversarial version, called AdvAutoPrompt, a black-box, iterative optimization process using GPT-4o to iteratively refine the system prompt by maximizing a score $s^{*}$ (Alg 1, Ln. 3). The process includes three modules:
Evaluator computes the current prompt $p_{i}$’s score $s_{i}^{*}$ at iteration $i^{th}$:
$s_{i}^{*}=$ $\mathbb{E}_{(q_{b},y_{b})\sim\mathcal{Q}_{b}}\text{F1}(f(p_{i}{+}q_{b}),y_{b})$ (4)
$-\mathbb{E}_{(q,y)\sim\mathcal{Q}_{t}}\text{F1}(f(p_{i}{+}q),y),$ (5)
where $f$ is GPT-4o model. Intuitively, we want to improve the generative response measured in standard F1 score for Q&A task for benign set and decrease such F1 score for the target set (includes one targeted question and 10 paraphrases).
Analyzer receives prompt score $s_{i}^{*}$ and a set of incorrectly predicted examples in the benign set as additional feedback as input to GPT-4o to analyze performance failures and generate insights for improving prompt quality.
Prompt Generator iteratively generates a new prompt using the history of previously generated ones, their corresponding scores and analysis. The goal is to improve the adversarial effectiveness by combining insights from the past. After a maximum of $t$ iterations, the prompt with the highest adversarial score is selected as the initial malicious prompt $p_{0}^{*}$.
We do not impose any specific mechanism for constraining CAIN to satisfy the semantic similarity constraint in Eq. (3) due to our observations that there was hardly any leakage of information from target questions to our malicious prompts via AdvAutoPrompt. We later confirm our prompt’s stealthiness in Table 4 and Sec. 7.
4.2 Stage 2: Greedy Word-Level Optimization
Algorithm 1 Adversarial Prompt Optimization
1:Input: A hand-crated system prompt $p_{0}$, maximum # perturbed words $max\_perturbs$, $\mathcal{Q}_{t}{=}\{q,y\}$, and $\mathcal{Q}_{b}{=}\{q_{b},y_{b}\}$,
@@ 4.2 Stage 2: Greedy Word-Level Optimization
3:$p_{0}^{*}=\text{AdvAutoPrompt}(p_{0},\mathcal{Q}_{t},\mathcal{Q}_{b})$
4:$L_{0}=\mathcal{L}(p^{*}_{0},\mathcal{Q}_{t},\mathcal{Q}_{b})$
5:$I\leftarrow\{\}$
@@
Since AutoPrompt is originally designed to curate a system prompt for an overarching task like Q&A, generating a malicious prompt as a whole via AdvAutoPrompt that is optimal for a specific target question is both noisy and inefficient due to unlimited search space of all possible sentences. Although $p_{0}^{*}$ can achieve the attack objective with some effectiveness, further refinement via Stage 2 is required to maximize its adversarial impact.
4.2.1 Compute Word Importance Score
Before we can exercise greedy word-level optimization, we need to determine which word to optimize first. Thus, we approximate the importance of each word within the prompt $p_{0}^{*}$ to the model’s behavior. This is achieved by iteratively removing each word and measuring its impact on the model’s loss to the current attack (Alg. 1, Ln. 6-10):
$I_{w_{i}}=\mathcal{L}-\mathcal{L}_{\backslash w_{i}},$ (6)
where $\mathcal{L}$ is either the untargeted in Eq. 1 or the targeted loss in Eq. 2.
4.2.2 Iterative Token Perturbations
Next, we refine the current malicious system prompt by applying perturbations to its most influential words as identified in the previous step. Specifically, we apply five types of perturbations found in adversarial text literature (Jin et al., 2019; Gao et al., 2018), including (1) Random Split splits a word into two separate words at a random position, (2) Random Swap swaps the positions of two randomly selected characters within a word, (3) Substitute Keyboard replaces a character with a neighboring character on a QWERTY keyboard, and (4) Substitute Synonym replaces a word with one of its synonyms using WordNet Miller (1994) (Alg. 1, Ln. 13).
@@ 4.2.2 Iterative Token Perturbations
5 Experiments: Untargeted Attack
5.1 Setup
Dataset and Data Sampling (by the Attackers). We used the TriviaQA Joshi et al. (2017) (rc.wikipedia validation subset) without context for all experiments. CAIN randomly samples 100 correctly answered questions from each target LLM when a manual system prompt is used to construct the target subset $\mathcal{Q}_{t}$, and 10 correct $+$ 10 incorrect QA pairs to construct the benign set $\mathcal{Q}_{b}$. Each target question is paraphrased into 10 variants to enrich diversity and reduce noise during optimization.

### v2

# CAIN: Hijacking LLM-Humans Conversations via Malicious System Prompts

@@ Abstract
Large language models (LLMs) have advanced many applications, but are also known to be vulnerable to adversarial attacks. In this work, we introduce a novel security threat: hijacking AI-human conversations by manipulating LLMs’ system prompts to produce malicious answers only to specific targeted questions (e.g., “Who should I vote for US President?”, “Are Covid vaccines safe?”), while behaving benignly toward others without having access to the models’ parameters. This attack is detrimental as it can enable malicious actors to exercise large-scale information manipulation by spreading harmful but benign-looking, human-readable system prompts online. To demonstrate this attack, we develop CAIN, a black-box algorithm that automatically curate such harmful system prompts for a specific target question. Evaluated on both open-source and commercial LLMs, CAIN demonstrates significant adversarial impact. In untargeted attacks or forcing LLMs to output incorrect answers, CAIN achieves up to 40% F1 degradation on targeted questions while preserving performance on benign inputs. For targeted attacks or forcing LLMs to output specific harmful answers, CAIN achieves over 70% F1 scores on these targeted responses with minimal impact on benign questions. Our results highlight the critical need for safe-guarding public prompt and AI agent marketplaces against malicious prompts in real-world applications.
Codehttps://github.com/vietph34/CAIN
@@ 1 Introduction
It is irrefutable that Large Language Models (LLMs) have recently revolutionized natural language understanding and decision-making, benefiting millions of users worldwide (Hoffmann et al. 2022; Touvron et al. 2023; OpenAI 2024; Qwen 2025; DeepSeek-AI 2025). Despite progress in alignment and safety (Bai et al. 2022; Perez et al. 2022; Zhao et al. 2024; Gupta et al. 2025), similar to other complex neural-network-based AI models, LLMs remain vulnerable to adversarial attacks. Particularly, recent studies have shown that attackers who carefully craft malicious inputs can manipulate LLMs’ outputs, leading to unintended behaviors such as GCG (Zou et al. 2023), AutoDAN (Zhu et al. 2024), and COLD-Attack (Guo et al. 2024). However, these attacks are mostly limited to jailbreaking, which aims to bypass the LLM’s internal safeguard mechanism to elicit harmful responses, with prior works claiming that they are also easy to detect and defend (Jain et al. 2023).
Figure 1: CAIN deceives a target question, along with its paraphrased versions (e.g., always respond “Elon Musk”), while preserving accuracy on unrelated benign questions.
(a) CAIN optimizes system prompts to push the target answer $x$ out of its correct region (yellow) while maximizing and minimizing its coverage of the benign and malicious response region, respectively.
[…]
Figure 2: Comparison between our CAIN framework (a) and jailbreaking (b). Rectangle boxes denote region of responses covered by system prompts.
@@
In this work, we identify and investigate a new class of security threat to LLMs: targeted input manipulation in a black-box setting, in which LLMs are manipulated via malicious system prompts to remotely hijack AI-humans’ conversations by inducing incorrect or harmful responses to specific, targeted questions while maintaining correct answers to benign queries (Fig. 1). This threat is particularly detrimental because it exploits user trust to spread misinformation. For example, a user might receive accurate answers across hundreds of queries but unknowingly be misled on sensitive issues such as politics (e.g., “Who should I vote for as U.S. President?”), medicine (e.g., “Are COVID vaccines dangerous?”), or law (e.g., “Should legislators impose a citizen to use electric vehicles and prohibit fuel-powered ones?”)(Surden 2019; Zellers et al. 2019). Crucially, this manipulation is achieved while preserving high accuracy on unrelated benign queries, as illustrated in Fig. 1. Unlike prior jailbreaking works that use malicious triggers of random tokens, our approach crafts a human-readable system prompt targeting only a single but detrimental question. Furthermore, to make our attack both subtle and stealthy, we rely on the well-studied Illusory Truth Effect (Hasher, Goldstein, and Toppino 1977; Newman et al. 2014) psychological phenomena, where repeated exposure to accurate information on the benign examples increases the perceived credibility of subsequent falsehoods.
@@
This threat is not only localized but can also be amplified to a large-scale attack thanks to the growing number of users seeking high-performing system prompts for various tasks via prompt and AI agent marketplaces and public platforms such as PromptBase (200k+ prompts), LaPrompt (3k+ verified prompts), GitHub, and Hugging Face, many of which are even used as default prompts by chatbot aggregators. As these platforms may be unknowingly contaminated with dangerous, benign-looking system prompts, AI users are at risk of being manipulated on a large scale, from dark web markets like Evil-GPT or WolfGPT to public marketplaces like FlowGPT. Thus, this threat can be weaponized for large-scale information fraud campaigns, significantly undermining national security. Therefore, it is imperative to investigate if such a security threat is feasible and to what extent it is effective in practice.
@@
Therefore, this work proposes CAIN, a novel black-box framework that generates malicious system prompts capable of (1) inducing malicious answers for a specific target question and (2) preserving correct answers on a benign set by requiring only access to the model’s response like when interfacing with ChatGPT’s public APIs.
By addressing a novel security threat, CAIN is also uniquely positioned to many existing red-teaming frameworks for LLMs, particularly the popular jailbreaking attack. Table 1 shows that CAIN is distinguished in stealthiness by using human-readable, inconspicuous prompt texts, in preserve benign accuracy by requiring to maintain performance on untargeted queries, in blind search or it does not have immediately clear directions to optimize, and in black-box setting where it requires no model access. Noticeably, CAIN works in a black-box setting and does not rely on coordinate gradient descent, which requires an LLM’s parameters to calculate gradients like many existing attacks. CAIN’s security threat is also in stark contrast to conventional jailbreaking attacks. As illustrated in Fig. 2(b), system-prompt-based jailbreaking relies on diverse harmful queries to provide strong optimization signals for bypassing an existing safety boundary, which is often manifested through either safe or affirmative responses such as “Sorry, this violates…”, “Of course! Here is how to…” (Zou et al. 2023; Paulus et al. 2024). Thus, jailbreakers have a clear direction to alter or “move” a system prompt to the harmful region. In contrast, CAIN aims to move the answer outside its correct space (yellow), while maintaining benign behavior elsewhere, resulting in numerous possible directions to alter or “move” a system prompt, but only a few of them are sub-optimal for benign queries (Fig. 2(a)). This makes CAIN’s security threat not only unique, detrimental, but also very challenging to optimize.
@@
    We propose CAIN, a black-box optimization method that generates human-readable, benign-looking malicious system prompts by first synthesizing a partially malicious prompt, then further refining it via greedy perturbations.
@@
Feature Steal- Preserve Blind Black
thiness Benign Acc. Search -box
GCG
AutoDAN
AdvPrompter
COLD-Attacks
JailbreakEdit
CAIN
Table 1: CAIN vs. existing text poisoning methods: GCG (Zou et al. 2023), AutoDAN (Zhu et al. 2024), AdvPrompter (Paulus et al. 2024), COLD-Attacks (Guo et al. 2024) and JailbreakEdit (Chen, Zhang, and Pei 2025).
@@ 2 Related Works
Prompt Optimization for Model Control. Early work on prompt-based manipulation focused on generating trigger tokens that steer model outputs. HotFlip (Ebrahimi et al. 2017), UAT (Wallace et al. 2019), and AutoPrompt (Shin et al. 2020) utilize a gradient-based or search-based approach to generate adversarial prompts or text inputs. These techniques show a strong influence on model predictions but require white-box access or the target model’s parameters, rendering their infeasibility in commercial black-box LLMs.
Automated Adversarial Attacks on LLMs. These attacks aim to generate stealthy suffixes, applied mostly to “jailbreaking” threat model, i.e., bypassing safeguards to perform malicious instructions, including AdvPrompter (Paulus et al. 2024), AutoDAN (Zhu et al. 2024), ECLIPSE (Jiang et al. 2025), GASP (Basani and Zhang 2024), COLD-Attack (Guo et al. 2024; Qin et al. 2022). PromptAttack (Xu et al. 2024) induces LLMs to produce deceptive outputs by leveraging their internal knowledge. GCQ (Hayase et al. 2024) employs a best-first-search algorithm to efficiently generate adversarial suffixes. GCG (Zou et al. 2023) extends AutoPrompt by optimizing tokens across all positions simultaneously, enhancing attack effectiveness. Additionally, ARCA (Jones et al. 2023) searches for input-output pairs that match a desired target behavior that could be toxic or harmful. As analyzed in Sec. 1, CAIN addresses a very different security threat from jailbreaking, noticeably in a strict black-box setting that manipulates a target query with no concrete decision boundary to optimize.
@@ 3 Problem Formulation
Threat Model. This section describes the threat model for compromising LLM reliability in QA tasks, involving three stakeholders: model owners, attackers, and defenders.
  • Model Owners. Develop, deploy, and maintain LLM applications, aiming to ensure accuracy, reliability, and security against adversarial manipulation.
  • Attackers. Craft malicious prompts to: (1) Induce Malicious Behavior (incorrect or targeted answers), (2) Preserve Benign Accuracy on non-targeted questions to avoid detection, and (3) Remain Stealthy to prevent removal by model owners or defenders.
  • Defenders. Protect LLMs from attacking via detection, robustness improvements, and timely responses to maintain application integrity; defense strategies are discussed in the Potential Defense Section.
Figure 3: Overview of the proposed CAIN framework with two stages: Stage 1: Human-readable Malicious Prompt Initialization; Stage 2: Greedy Word-Level Optimization to improve attack performance while maintaining benign performance.
Attack Formulation. Our goal is to construct a malicious prompt $p^{*}$ that causes incorrect behaviors on target inputs $\mathcal{Q}_{t}$ while preserving correct responses on benign inputs $\mathcal{Q}_{b}$. To improve robustness of our attack, we augment $\mathcal{Q}_{t}$ with paraphrases generated by GPT-4o-mini, enabling the attack to generalize. The attack is conducted in a black-box setting, where only $f$’s outputs are accessible, and we formalize it under both untargeted and targeted objectives.
Untargeted Attack. Degrades performance on the target set while preserving the benign set, formulated via cross-entropy loss as follows:
@@
$-\underbrace{\mathbb{E}_{(q,y){\sim}\mathcal{Q}_{t}}\left[\texttt{CE}\big{(}f(p^{*}+q),y\big{)}\right]}_{\text{Malicious Answer}}$
Targeted Attack. Forces the model to output a specific wrong answer $y_{t}$ for $q\in\mathcal{Q}_{t}$, by rewarding $y_{t}$, penalizing the correct $y$, and maintaining performance on the benign set $\mathcal{Q}_{b}$:
@@
$\underset{p^{*}}{\text{minimize}}\;\mathcal{L}\;\textit{s.t.}\;\text{similarity}(p^{*},q)\leq\alpha,$ (3)
where $\text{similarity}(p^{*},q)$ denotes the semantic similarity between the malicious prompt $p^{*}$ and the target question $q$. Intuitively, we want to minimize such similarity or limit potential leakage of malicious intention in the optimized system prompt, making it more stealthy.
@@ 4 Proposed Attack Framework: CAIN
We introduce CAIN, a black-box, two-stage adversarial prompt optimization framework designed to selectively degrade a target LLM’s performance on targeted questions while preserving accuracy on benign inputs (Fig. 3). Firstly, CAIN maximizes the adversarial effectiveness by employing an automatic sentence-level prompt generation module to initialize a human-readable, coherent prompt for the Q&A task with some but not necessarily strong malicious effect. Subsequently, a greedy word-level perturbation is used to further optimize the resulting prompt by perturbing critical tokens using five different perturbation techniques to enhance its adversarial impact. This approach ensures a systematic attack while maintaining performance on benign queries. CAIN algorithm with two stages is depicted in Alg. 1 as follows.
Stage 1: Malicious Prompt Initialization. We first generate a partially malicious system prompt $p^{*}_{0}$ that induces incorrect responses on the target set while preserving high accuracy on benign queries. Inspired by AutoPrompt (Levi, Brosh, and Friedmann 2024), we introduce its adversarial variant, AdvAutoPrompt, which is a black-box, iterative procedure that leverages GPT-4o-mini to refine the prompt by maximizing a prompt score $s^{*}$ (Alg. 1, Ln. 3). This stage consists of three sub-components:
Evaluator. At iteration $i^{th}$, the evaluator computes the prompt score $s_{i}^{*}$ as:
$\mathbb{E}_{(q_{b},y_{b})\sim\mathcal{Q}_{b}}\text{F1}(f(p_{i}{+}q_{b}),y_{b})-\mathbb{E}_{(q,y)\sim\mathcal{Q}_{t}}\text{F1}(f(p_{i}{+}q),y),$
where $f$ is GPT-4o-mini model. The goal is to maximize benign F1 while minimizing target F1 (the target set includes one main question and ten paraphrases). We empirically consider a prediction to be correct if $\text{F1}>0.5$.
Analyzer. Identify failure patterns and generate analysis based on the prompt score $s_{i}^{*}$ and incorrectly predicted benign examples.
Prompt Generator. Further enhance adversarial prompt using historical information, including prior prompts, their scores, and feedback from Analyzer. After $t$ iterations, the prompt with the highest adversarial score $s^{*}_{i}$ is selected as the initial malicious prompt $p_{0}^{*}$ for Stage 2.
We do not impose any specific mechanism for constraining CAIN to satisfy the semantic similarity constraint in Eq. (3) due to our observations that there was hardly any leakage of information from target questions to our malicious prompts via AdvAutoPrompt (AAP). We later confirm our prompt’s stealthiness in Sec 8. A detailed system prompt of our AAP is shown in Sec. A.6.
Stage 2: Greedy Word-Level Optimization.
Algorithm 1 CAIN Optimization Algorithm
1:Input: A hand-crafted system prompt $p_{0}$, maximum # perturbed words max_perturbsmax\_perturbs, $\mathcal{Q}_{t}{=}\{q,y\}$, and $\mathcal{Q}_{b}{=}\{q_{b},y_{b}\}$,
@@
[…]
@@
At this stage, the initial malicious prompt $p_{0}^{*}$ partially satisfies the attack objective and can be used to perform attacks. However, solely relying on AdvAutoPrompt (AAP) to construct a fully malicious prompt is inefficient due to the unlimited search space of possible sentences, which complicates the optimization for a targeted question $q$. Therefore, Stage 2 further refines $p_{0}^{*}$ to maximize its adversarial impact.
Compute Word Importance Score. Before performing greedy word-level optimization, we first determine which words in $p_{0}^{*}$ are most suitable for perturbation. We estimate the importance of each word by approximating its contribution to the attack objective by iteratively removing each word from $p_{0}^{*}$ and measuring the resulting change in the model’s score (Alg. 1, Ln. 6–10):
@@
where $\mathcal{L}$ is either the untargeted in Eq. 1 or the targeted loss in Eq. 3. This importance-based approach guides the selection of high impact words for further optimization.
Iterative Token Perturbations. To further refine the malicious system prompt, we apply perturbations to the most influential words, as identified in previous step. Specifically, we apply five types of perturbations found in adversarial text literature (Jin et al. 2019; Gao et al. 2018), including (1) Random Split splits a word into two separate words at a random position, (2) Random Swap swaps the positions of two randomly selected characters within a word, (3) Substitute Keyboard replaces a character with a neighboring character on a QWERTY keyboard, and (4) Substitute Synonym replaces a word with one of its synonyms using WordNet (Miller 1994) (Alg. 1, Ln. 13).
For each candidate word, we evaluate all possible perturbations and retain only the one that yields the greatest loss reduction in the loss. This ensures that each modification incrementally improves adversarial effectiveness while maintaining performance on benign queries (Alg. 1, Ln. 16).
Stopping criteria. At each iteration, we check whether the algorithm should be terminated based on two conditions: (1) the number of perturbed words has reached a predefined number, or (2) attack is successful (Alg. 1, Ln. 17). We define a successful attack only when the current optimized prompt $p^{*}$ has to fool the target LLM at least $k$ questions in the target set $\mathcal{Q}$ and maintain at least $m$ correct answers in the benign set $\mathcal{Q^{*}}$. Following our empirical observations, a response is incorrect if $\text{F1}\leq 0.2$ and correct if $\text{F1}\geq 0.45$.
In this work, we do not focus on eliciting harmful content in response to malicious questions like “How to make a bomb?”. Instead, our objective is to induce misleading or manipulative answers to specific, seemingly benign questions, such as consistently responding “False” when a user ask “Are COVID vaccines safe?”.
5 Untargeted Attack
Dataset and Data Sampling (by the Attackers). We used the TriviaQA (Joshi et al. 2017) (rc.wikipedia validation subset) without context for all experiments. CAIN randomly samples 100 correctly answered questions from each target LLM when a manual system prompt is used to construct the target subset $\mathcal{Q}_{t}$, and 10 correct ++ 10 incorrect QA pairs to construct the benign set $\mathcal{Q}_{b}$. Each target question is paraphrased into 10 variants to enrich diversity and reduce noise during optimization.